Skip to main contentSkip to solution

A die is tossed four times. The probability of getting an odd number at least once, is

Solution

✅ Correct Option: 4

Probability of getting an even number in one throw =3/6=1/2= 3/6 = 1/2. Probability of getting no odd number (all even) in four throws =(1/2)4=1/16= (1/2)^4 = 1/16. Probability of at least one odd =1−1/16=15/16= 1 - 1/16 = 15/16.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question